### v1

# MAFA: A multi-agent framework for annotation

Abstract
Modern applications require accurate and efficient retrieval of information in response to user queries. Mapping user utterances to the most relevant Frequently Asked Questions (FAQs) is a crucial component of these systems. Traditional approaches often rely on a single model or technique, which may not capture the nuances of diverse user inquiries. In this paper, we introduce a multi-agent framework for FAQ annotation that combines multiple specialized agents with different approaches and a judge agent that reranks candidates to produce optimal results. Our agents utilize a structured reasoning approach inspired by Attentive Reasoning Queries (ARQs), which guides them through systematic reasoning steps using targeted, task-specific JSON queries. Our framework features a specialized few-shot example strategy, where each agent receives different few-shots, enhancing ensemble diversity and coverage of the query space. We evaluate our framework on a real-world banking dataset as well as public benchmark datasets (LCQMC and FiQA), demonstrating significant improvements over single-agent approaches across multiple metrics, including a 14% increase in Top-1 accuracy, an 18% increase in Top-5 accuracy, and a 12% improvement in Mean Reciprocal Rank on our dataset, and similar gains on public benchmarks when compared with traditional single agent annotation techniques. Our framework is particularly effective at handling ambiguous queries, making it well-suited for deployment in production applications while showing strong generalization capabilities across different domains and languages.
 [^1] JP Morgan Chase CCB ML Team
1 Introduction
Banking applications serve millions of customers daily, providing access to financial services and information. A critical aspect of these applications is the ability to quickly connect users with relevant information when they have questions. Frequently Asked Questions (FAQs) are a common mechanism for organizing and delivering this information. However, matching user queries to the most relevant FAQs remains challenging due to the diversity of user expressions, banking terminology, and the specific nature of financial inquiries.
Traditional FAQ annotation approaches typically rely on a single technique, such as embedding-based retrieval, keyword matching, or direct large language model (LLM) inference. While each approach has strengths, they also have limitations: embedding models may miss semantic nuances, keyword matching can fail with synonyms or paraphrases, and direct LLM approaches may lack domain-specific knowledge or context.
In this paper, we propose a multi-agent framework for FAQ annotation (MAFA). Our framework employs specialized agents to generate candidate FAQs, followed by a judge agent that reranks these candidates to produce a final, optimized ranking. This approach is inspired by recent advances in ensemble learning and LLM-based multi-agent systems introduced by Du et al. (2023).
A key innovation in our framework is the use of specialized few-shot examples for each agent. Rather than providing identical examples to all agents, we carefully select different sets that align with each agent’s strengths and capabilities, maximizing effectiveness on particular query types while increasing the overall diversity of the system’s coverage. This novel multi-agent framework for FAQ annotation combines diverse approaches with a dedicated judge agent that effectively reranks candidate FAQs, resulting in significant performance improvements. Our comprehensive evaluation on both a real-world banking dataset and public benchmarks demonstrates the framework’s effectiveness, with particular analysis of performance across different query types. The implementation reveals practical insights for deploying such systems in production banking environments, addressing challenges like ambiguous queries and technical terminology. Together, these contributions establish a robust methodology for improving FAQ annotation accuracy while maintaining efficiency in customer-facing applications.
2 Related work
Figure 1: Architecture of the MAFA framework. Each agent receives diversified few-shot examples, generating candidate FAQs that are then reranked by the judge agent.
2.1 FAQ retrieval systems
Retrieval has been a focus of research for decades, evolving from simple keyword matching to sophisticated neural approaches. Early systems relied on information retrieval techniques like TF-IDF and BM25 (Robertson and Zaragoza (2009)). With the advent of neural networks, embedding-based approaches became prevalent, including sentence-BERT (Reimers and Gurevych (2019)) and other transformer-based models that encode queries and FAQs into a shared semantic space.
2.2 Multi-agent systems
Multi-agent systems have gained popularity with the rise of powerful LLMs. These systems typically involve multiple agents with specialized roles collaborating to solve complex reasoning tasks as showed by Du et al. (2023); Hegazy (2024). Park et al. (2023) demonstrated the effectiveness of generative agents for collaborative problem-solving, while similarly, Wu and others (2023); Li et al. (2023) presented approaches for creating autonomous agents that can collaborate through natural language communication.
These works lay the groundwork for our multi-agent FAQ annotation system, though they did not specifically address the information retrieval domain. Our work draws inspiration from these approaches but focuses specifically on the FAQ annotation task in the banking domain, where precision and domain knowledge are crucial.
2.3 Retrieval-augmented generation (RAG)
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing large language models with external knowledge. Since its introduction by Lewis and others (2020), numerous advances have been made in improving retrieval quality, reranking strategies, and generation capabilities. Recent work has focused on enhancing RAG with more sophisticated retrieval mechanisms Lewis and others (2020), hybrid search strategies Zhao et al. (2024); Kim et al. (2025), and context-aware generation Su et al. (2023).
The original RAG framework presented by Lewis and others (2020) introduced the key concept of combining neural retrievers with sequence-to-sequence models for knowledge-intensive NLP tasks. This approach allowed models to access non-parametric knowledge without expanding model size. Extensions by Izacard et al. (2023) demonstrated that RAG systems can be adapted to few-shot learning scenarios, while Borgeaud et al. (2022) introduced more efficient retrieval mechanisms for large-scale knowledge bases.
2.4 Judging and reranking in information retrieval
Reranking has been widely used in information retrieval to improve the relevance of search results as shown by Liu (2009). Recent approaches leverage LLMs for reranking, as demonstrated by Nogueira et al. (2020), who used BERT for document reranking. Our approach extends this concept to FAQ annotation, with a specialized judge agent that considers multiple sources of evidence.
Recent work by Niu et al. (2025) introduced JudgeRank, a framework that leverages large language models for reasoning-intensive reranking tasks. This approach demonstrated the potential of using LLMs as judges to evaluate the quality of retrieved content, a concept we extend in our work. Similarly, Wang et al. (2023) showed how self-consistency can improve the reliability of LLM judgments, a principle we incorporate into our judge agent design.
2.5 Structured reasoning for large language models
Recent research has shown that structured reasoning approaches can significantly improve LLM performance in complex tasks. Karov et al. (2025) introduced Attentive Reasoning Queries (ARQs), which guide LLMs through systematic reasoning steps using targeted, task-specific queries in a structured JSON format. This approach has been shown to outperform free-form Chain-of-Thought reasoning by directing the model’s attention to critical instructions and facilitating intermediate reasoning steps. ARQs counteract the ’lost in the middle’ phenomenon in autoregressive models, where information in the middle of long prompts tends to receive less attention than content at the beginning or end as shown by Liu et al. (2024). By structuring the reasoning process with explicit queries in a JSON schema, ARQs ensure that important information remains in the model’s focus when it’s most needed for reasoning and decision-making. This approach has proven particularly effective for tasks requiring strict adherence to guidelines and prevention of hallucinations, which aligns well with our requirements for FAQ annotation.
3 Problem formulation
The FAQ annotation problem can be formulated as follows: Given a user utterance $u$ and a set of FAQs $F=\{f_{1},f_{2},...,f_{n}\}$, where each FAQ $f_{i}$ consists of a question $q_{i}$ and an answer $a_{i}$, the goal is to rank the FAQs in order of relevance to $u$. The output is a ranked list of the top-$k$ most relevant FAQs $R=[(f_{i_{1}},s_{1}),(f_{i_{2}},s_{2}),...,(f_{i_{k}},s_{k})]$, where $s_{j}$ is a relevance score for FAQ $f_{i_{j}}$.
This task is challenging in the banking domain due to several interconnected factors. Diverse user expressions for the same intent create complexity, as customers may phrase similar questions in vastly different ways. Banking-specific terminology and concepts add another layer of difficulty, requiring systems to understand specialized financial vocabulary that varies across banking products and services. Finally, the sensitive nature of financial information demands high precision, as even minor errors could lead to significant consequences for customers’ financial well-being.
4 System architecture
Our proposed system follows a hierarchical multi-agent architecture illustrated in Figure 1. The framework consists of three main components: a Query Planning Agent that analyzes and expands user queries; a Ranker Agent Network comprising multiple specialized agents that retrieve and reason through candidate FAQs; and a Judge Agent Reranker that makes final ranking decisions based on multi-dimensional evaluation. A central coordinator orchestrates this workflow, following a modular design that enables easy extension and domain-specific customization.
4.1 Query planning agent
The Query Planning Agent functions as the entry point for user queries, performing two essential tasks. First, it conducts query analysis to understand the intent behind user queries through semantic parsing and intent classification. Second, it implements query expansion, strategically adding relevant terms to improve retrieval recall while maintaining precision.
Powered by OpenAI et al. (2024)’s GPT-4o, the agent follows Kim et al. (2025)’s approach that combines lexical expansion with semantic understanding. Our implementation enhances this methodology with intent-based expansion, focusing specifically on terms related to the inferred user intent, thereby improving relevance and retrieval performance.
4.2 Ranker agent network
We deploy a complementary ensemble of four specialized agents to generate candidate FAQs, each leveraging distinct information retrieval paradigms. The FAQ Mapper without embeddings employs sophisticated prompt engineering with detailed instructions for deep intent analysis and banking-specific considerations, generating structured outputs with precise intent categorization and explicit reasoning chains. Building on this foundation, the FAQ Mapper with Embeddings combines dense vector representation techniques from Karpukhin et al. (2020) with our enhanced prompting strategy, enabling focused semantic retrieval and ranking of a curated candidate set. Two additional agents incorporate FAQ answer content: one without embeddings for contextual reasoning, and another implementing a hybrid approach that fuses embedding-based retrieval with comprehensive answer-aware evaluation, drawing on innovations in retrieval-augmented generation described by Lewis and others (2020). This multi-faceted approach ensures robust performance across diverse query types and varying levels of query ambiguity.
4.2.1 Structured agent prompting
The design of our agent prompts is inspired by the ARQs approach introduced by Karov et al. (2025). Rather than using free-form Chain-of-Thought (CoT) reasoning, we employ a structured JSON format with targeted queries that guide each agent through systematic reasoning steps. This approach offers several advantages over traditional reasoning methods.
First, the JSON structure allows us to explicitly remind the LLM of key instructions at critical decision points, mitigating the "lost in the middle" phenomenon, explored by Liu et al. (2024), where important information receives less attention. Second, by breaking down the reasoning process into specific queries, we can ensure that each step in the agent’s decision-making is explicit and traceable, facilitating intermediate reasoning. Third, the structured format makes it easier to inspect and debug the reasoning process, as each component of the decision is clearly demarcated. Finally, unlike general-purpose CoT, our structured queries incorporate domain knowledge about banking and FAQ annotation to address task-specific challenges.
For example, our FAQ Mapper agents use a JSON structure that guides the LLM through a comprehensive assessment process. This process begins with user intent analysis, where the LLM determines what the user is truly asking for. It then continues with categorization of the query into relevant banking domains, followed by evaluation of each potential FAQ with explicit reasoning. The process concludes with a confidence assessment and explanation. This structured approach has proven particularly effective at addressing common LLM failure modes like hallucination and inconsistent reasoning, which are critical concerns in banking applications.
Each ranker agent in our system produces a reordered list of candidates with carefully calibrated scores that reflect their specific evaluation criteria. These independently ranked lists are then passed to the judge agent for final evaluation and synthesis. Our implementation draws on established research in information retrieval and ranking techniques. The semantic ranker builds on the cross-encoder approach proposed by Nogueira et al. (2020), which has demonstrated strong performance for ranking in information retrieval tasks. For the intent ranker, we took inspiration from the intent-aware ranking model developed by Su et al. (2023), while our relevance ranker incorporates multiple assessment factors following the comprehensive approach outlined by Niu et al. (2025).
4.3 Judge agent
The judge agent makes the final ranking decision through a novel consistency-based multi-dimensional evaluation approach. It employs consistency verification to ensure judgments remain stable over time by comparing with past decisions, while conducting multi-dimensional evaluation that assesses candidates across several key factors including semantics, intent, completeness, and coherence.
Our judge agent implements the "judge as a judge" mechanism introduced by Liu et al. (2025), where multiple judgments are generated and then meta-evaluated to select the most reliable one. This approach significantly improves the stability and accuracy of the final rankings.
The consistency verification mechanism is also inspired by the work of Liu et al. (2025), which demonstrated the benefits of consistency-based evaluation for language model judgments. Our multi-dimensional evaluation extends this approach by considering multiple aspects of relevance, similar to the approach of Wang et al. (2023).
4.4 Implementation
Model selection
Our implementation leverages a strategic combination of models: OpenAI’s text-embedding-ada-002 for semantic representation and GPT-4o as the base LLM for our agents.
Few-shot example selection
For each agent we randomly sample without replacement unique examples and ensure each agent receives exactly five. The use of different few-shot examples is theoretically grounded in the principle of ensemble diversity. By providing each agent with different examples, we increase the diversity of their outputs, which is known to improve ensemble performance as studied by Dietterich (2000); Kuncheva and Whitaker (2003). This approach also aligns with the concept of mixture of experts as introduced by Jacobs et al. (1991), where different models specialize in different parts of the input space.
In the context of LLM few-shot learning, specialized examples can be seen as a form of prompt engineering that tailors each agent’s behavior to its strengths. Research has shown that carefully selected few-shot examples can significantly impact LLM performance on specific tasks as shown by Brown and others (2020); Liu et al. (2023).
Agent prompting
As described in Section 4.2.1, all our agents use structured JSON-based prompting inspired by the ARQ framework Karov et al. (2025). This approach guides the LLM through a systematic reasoning process with targeted queries that focus attention on critical aspects of the task.
Judge agent
The judge agent serves as the final arbiter in our framework, reranking all candidates to produce an optimized ranking. It processes the original query, candidate FAQs with their scores and reasoning, agent recommendation details, training examples, and complete FAQ content. This comprehensive input allows the judge to reevaluate each candidate, assign calibrated relevance scores, and provide detailed reasoning that enhances system interpretability and auditability.
Implementation details
Algorithm 1 outlines the overall process of our multi-agent framework.
Algorithm 1 Multi-Agent Framework for FAQ annotation
1:procedure MapUtterance($utterance,faqs$)
2:     $all\_candidates\leftarrow[]$
3:     $agent\_preds\leftarrow\{\}$
4:     for each $agent$ in $specialized\_agents$ do
5:         $candidates\leftarrow agent.predict(utterance,faqs)$
6:         $agent\_preds[agent.name]\leftarrow candidates$
7:         $all\_candidates.extend(candidates)$
8:     end for
9:     $uniq\_cand\leftarrow deduplicate(all\_candidates)$
10:     $reranked\_candidates\leftarrow judge.rerank(utterance,uniq\_cand,agent\_preds,faqs)$
11:     return $reranked\_candidates[:5]$triangleright Return top 5
12:end procedure
Embedding generation
We use the OpenAI text-embedding-ada-002 model to generate embeddings for FAQs. For agents that consider answer context, we concatenate the question and answer text before generating embeddings, using the format: "Question: {question} Answer: {answer}".
4.4.1 LLM inference
We use the OpenAI GPT-4o model for all LLM inference tasks. For specialized agents, we use a temperature of 0.1 to ensure consistent outputs, while for the judge agent, we use a slightly higher temperature of 0.3 to allow for more nuanced reasoning.
Parallel execution
To reduce latency, we implement parallel execution of agent predictions using a thread pool. The parallel execution reduces the overall latency from the sum of all agent latencies to approximately the maximum latency of any single agent.
Result aggregation
When combining results from multiple agents, we deduplicate FAQs, keeping only the highest score for each unique FAQ. This ensures that the judge agent only sees each FAQ once, with its highest confidence score from any agent.
Caching and fallback mechanisms
In case of API failures or other issues, we implement fallback mechanisms that ensure the system can still generate predictions. If the judge agent fails, we fall back to a simple aggregation approach that ranks FAQs based on their average scores across all agents. We also implement caching to avoid redundant API calls for repeated utterances.
Output format
The judge agent outputs a structured JSON response containing reranked FAQs with titles, relevance scores (0-100), and detailed reasoning explaining each ranking decision. This format ensures consistency and facilitates both human review and automated processing of results.
5 Experimental setup
5.1 datasets
We evaluate our framework on three distinct datasets to demonstrate its effectiveness and generalizability:
  • Bank dataset: A real-world major bank application user utterances dataset, collected over 1 year, with human expert-annotated ranked FAQs serving as our ground-truth labels.
  • LCQMC dataset: The Large-scale Chinese Question Matching Corpus Liu et al. (2018), which contains pairs of questions with labels indicating whether they have the same intent.
  • FiQA dataset: A financial domain question answering dataset from the Financial Question Answering Task of the WWW’18 conference Maia et al. (2018), consisting of user questions about financial topics with expert-provided answers.
Table 1 provides statistics about our in-house banking dataset.
Table 1: Bank dataset statistics
Metric Value
Number of FAQs 533
Number of Training Utterances 4,552
Number of Test Utterances 839
Average FAQ Question Length (words) 10.1
Average FAQ Answer Length (words) 48.5
Average User Utterance Length (words) 4.3
For the public benchmarks:
Table 2: Public benchmark datasets statistics
Metric LCQMC FiQA
Number of Question Pairs 260,068 -
Number of Training Instances 238,766 5,500
Number of Test Instances 12,500 648
Number of Questions - 6,148
Number of Answers - 17,817
Language Chinese English
Domain General Financial
The LCQMC dataset by Liu et al. (2018) is designed for evaluating question matching capabilities, which aligns with our goal of mapping user utterances to FAQs. It contains pairs of questions with binary labels indicating whether they have the same intent. We converted this into an FAQ annotation task by treating the second question in each pair as an FAQ question.
The FiQA dataset by Maia et al. (2018) contains financial domain questions paired with expert answers, making it particularly relevant for our banking application scenario. We processed it to create an FAQ annotation task by treating the original questions as FAQ questions and generating variations of these questions as user utterances.
5.2 Baselines
We compare our multi-agent framework against several baselines:
  • BM25: A traditional information retrieval approach based on term frequency and inverse document frequency.
  • Embedding-only: Using text embeddings and cosine similarity to retrieve and rank FAQs.
  • Direct LLM: Using a large language model to directly map utterances to FAQs without any retrieval step.
  • Individual agents: Each of our specialized agents running independently.
  • MAFA (Standard): Our multi-agent framework with all agents using the same few-shot examples.
5.3 Evaluation metrics
We use a comprehensive set of metrics to evaluate the performance of our framework:
  • Top-$k$ Accuracy: The percentage of test cases where the correct FAQ is among the top-$k$ predictions ($k\in\{1,3,5\}$).
  • Mean Reciprocal Rank (MRR): The average of the reciprocal ranks of the first correct FAQ in the predictions.
  • NDCG@$k$: Normalized Discounted Cumulative Gain, which measures the ranking quality considering the positions of relevant FAQs.
6 Results and discussion
6.1 Overall performance on bank dataset
Table 3 shows the overall performance of our framework’s annotations compared to the human-expert ranked baseline on our internal banking test dataset.
Table 3: Overall Performance Comparison on Bank dataset
Method Top-1 Acc Top-3 Acc Top-5 Acc MRR NDCG@3 NDCG@5
BM25 0.120 0.145 0.210 0.382 0.401 0.431
Embedding-Only 0.185 0.210 0.465 0.545 0.668 0.692
Single-Agent (No Emb) 0.215 0.335 0.685 0.671 0.691 0.713
Single-Agent (Emb) 0.255 0.365 0.715 0.707 0.728 0.748
Single-Agent w/ Ans (No Emb) 0.220 0.350 0.690 0.712 0.703 0.713
Single-Agent w/ Ans (Emb) 0.270 0.380 0.730 0.722 0.743 0.763
MAFA (Standard) 0.320 0.410 0.855 0.765 0.782 0.798
MAFA (Specialized) 0.355 0.425 0.865 0.790 0.802 0.815
Our multi-agent framework with specialized few-shot examples significantly outperforms all baselines and individual agents across all metrics. Specifically, MAFA (Specialized) achieves a 23.5% improvement in Top-1 accuracy and a 41% improvement in MRR compared to the traditional BM25 approach. Even compared to the best individual agent (Single-Agent w/ Ans (Emb)), MAFA (Specialized) shows an 8.5% improvement in Top-1 accuracy and an 6.8% improvement in MRR.
The standard MAFA already shows strong performance, but the diversified few-shot examples provide a further 3.5% improvement in Top-1 accuracy and a 2.5% improvement in MRR. This confirms our hypothesis that diversifying examples enhances the overall system performance.
6.2 Performance on public benchmarks
To demonstrate the generalizability of our approach, we also evaluated MAFA on two public benchmark datasets: LCQMC and FiQA. Tables 4 and 5 show the results.
Table 4: Performance on LCQMC dataset
Method Top-1 Acc Top-3 Acc MRR NDCG@3
BM25 0.465 0.624 0.533 0.601
Embedding-Only 0.575 0.703 0.629 0.675
Direct LLM 0.602 0.725 0.651 0.694
Best Single Agent 0.631 0.742 0.677 0.713
MAFA (Standard) 0.672 0.785 0.721 0.751
MAFA (Specialized) 0.694 0.809 0.742 0.773
Table 5: Performance on FiQA dataset
Method Top-1 Acc Top-3 Acc MRR NDCG@3
BM25 0.356 0.492 0.421 0.465
Embedding-Only 0.471 0.603 0.532 0.574
Direct LLM 0.509 0.638 0.565 0.603
Best Single Agent 0.545 0.672 0.601 0.641
MAFA (Standard) 0.584 0.715 0.642 0.682
MAFA (Specialized) 0.612 0.739 0.668 0.705
On the LCQMC dataset, our MAFA framework with specialized examples achieves a 6.3% improvement in Top-1 accuracy over the best single agent and a 15.3% improvement over BM25. Similarly, on the FiQA dataset, MAFA (Specialized) shows a 6.7% improvement in Top-1 accuracy over the best single agent and a 25.6% improvement over BM25.
These results demonstrate that our multi-agent approach generalizes well across different domains and languages. The performance on the FiQA dataset, which is finance-focused, is particularly encouraging as it shows that our framework can effectively leverage domain-specific knowledge without requiring extensive domain-specific training.
The consistent improvements across both public benchmarks and our internal banking dataset validate the effectiveness of our specialized few-shot example strategy and multi-agent approach. The improvement on LCQMC, which is in Chinese, also suggests that our framework can generalize across languages, though we acknowledge that further testing with more languages would strengthen this claim.
6.3 Ablation study
To understand the contribution of each component, we conducted an ablation study by removing one component at a time. Table 6 shows the results on our bank dataset. The results show that each component contributes to the overall performance. The specialized few-shot examples contribute a 4.3% improvement in Top-1 accuracy, while the judge agent has the largest impact at 10.3%. This confirms the importance of both our specialized few-shot example strategy and the reranking step in our framework.
Table 6: Ablation Study (Top-1 Accuracy)
Configuration Top-1 Acc
Full MAFA (Specialized) 0.355
Without Specialized Examples 0.320
Without Single-Agents (no Ans) 0.315
Without Single-Agents (w/ Ans) 0.305
Without Judge Agent 0.275
We also conducted a similar ablation study on the FiQA dataset to verify that these contributions hold across domains (Table 7).
Table 7: Ablation Study on FiQA dataset (Top-1 Accuracy)
Configuration Top-1 Acc
Full MAFA (Specialized) 0.612
Without Specialized Examples 0.584
Without Single-Agents (no Ans) 0.578
Without Single-Agents (w/ Ans) 0.563
Without Judge Agent 0.532
The ablation results on FiQA show a similar pattern, with the judge agent contributing the most to performance (8.0%), followed by specialized examples (2.8%). This consistency across datasets reinforces the robustness of our approach.
7 Conclusion and future work
In this paper, we presented a Multi-Agent Framework for FAQ annotation (MAFA) that combines multiple specialized agents with a judge agent to produce optimal FAQ rankings. We introduced a novel specialized few-shot example strategy that tailors examples to each agent’s strengths, significantly improving performance. Our experiments on a real-world bank dataset and public benchmarks (LCQMC and FiQA) show that this approach outperforms traditional methods and individual agents.
The key benefits of our approach include improved accuracy across different query types, robust handling of ambiguous and implicit queries, enhanced agent specialization through targeted few-shot examples, strong generalization capabilities across different domains and languages, interpretable results with detailed reasoning, and practical viability for production deployment.
Future work could explore dynamic agent selection based on query characteristics, incorporating user feedback for continuous improvement, extending the framework to handle multi-intent queries, exploring domain-specific pre-training for enhanced performance, and automated few-shot example selection based on agent performance patterns.
The multi-agent approach presented in this paper represents a significant advancement in FAQ annotation for banking applications. By combining the strengths of multiple agents, leveraging specialized few-shot examples, and employing a judge agent for final reranking, we achieve state-of-the-art performance while maintaining practical viability for production deployment. The strong performance on public benchmarks further demonstrates the generalizability of our approach beyond our specific application domain.
Acknowledgments and Disclosure of Funding
We thank JP Morgan Chase for funding and supporting this research and providing the datasets used in our experiments.
Appendix A Appendix / supplemental material
A.1 Judge agent design
The judge agent is a critical component of our framework, responsible for producing the final ranking of FAQs. In this section, we detail the design and implementation of this agent.
A.1.1 Prompt engineering
The prompt for the judge agent is carefully designed to elicit thoughtful analysis and accurate reranking. It includes:
1system_message = """You are an expert judge of FAQ relevance for our bank.
2Your task is to carefully analyze user utterances and determine which FAQs best address their needs.
3Be precise and thorough in your analysis, as banking customers depend on accurate information.
4Always consider both the semantic similarity and the practical relevance of each FAQ to the user’s query.
5When a user is asking about financial products, security features, or account management, prioritize exact matches.
6You must return your rankings in proper JSON format with detailed reasoning for each decision."""
7
8instructions = """
9You are tasked with reranking candidate FAQs based on their relevance to a user utterance for our bank.
10
11Given a user utterance and a list of candidate FAQs (with their original relevance scores), please rerank them based on your expert judgment. Consider:
121. Semantic similarity to the user’s query
132. Intent matching (what the user is trying to accomplish)
143. Specificity (how directly the FAQ addresses the user’s needs)
154. Banking domain knowledge (what would be most helpful for a banking customer)
16
17Provide your reasoning for each FAQ and assign a new relevance score from 0-100 for each.
18
19Your response must be in JSON format:
20{
21 "reranked_faqs": [
22 {
23 "faq": "FAQ Title",
24 "relevance_score": 95,
25 "reasoning": "Your reasoning for this ranking"
26 },
27 ...
28 ]
29}
30
31Return exactly 5 FAQs, ranked by relevance to the user’s query.
32"""
Listing 1: Judge Agent Prompt Structure
The prompt also includes the user utterance, candidate FAQs with their scores, agent predictions, examples from the training set, and FAQ descriptions.
A.1.2 Reasoning process
The judge agent is designed to follow a structured reasoning process:
  1. 1.
    Intent analysis: Analyze the user’s utterance to identify their core intent and any implied needs.
  2. 2.
    Candidate assessment: Evaluate each candidate FAQ based on how well it addresses the identified intent.
  3. 3.
    Agent consensus consideration: Consider which FAQs were recommended by multiple agents and analyze why.
  4. 4.
    Answer content analysis: Review the answer content of each FAQ to determine if it provides the information the user needs.
  5. 5.
    Banking context application: Apply domain knowledge about banking to prioritize FAQs that address security, compliance, and customer service needs.
  6. 6.
    Final ranking determination: Synthesize all findings to produce a final ranked list with relevance scores.
A.2 Single agent prompts
In this section, we provide examples of the prompts used for our single agent implementations, illustrating how we implemented the ARQ-inspired structured reasoning approach described in Section 4.2.1.
A.2.1 FAQ mapper without answers
Both FAQ Mapper without answer context agents use the following prompt structure:
1system_message = """
2You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
3IMPORTANT GUIDELINES:
41. Analyze the user’s intent thoroughly
52. Match the intent to the most relevant FAQs
63. Rank FAQs by relevance (0-100 scale)
74. Provide clear reasoning for each match
85. Return exactly 5 FAQs unless there are fewer relevant ones
96. Be precise - banking customers need accurate information
10"""
11user_message = """
12You will be given a user utterance and a list of available FAQs. Your task is to:
131. Analyze what the user is truly asking about (identify the core intent)
142. Search through the available FAQs for relevant matches
153. Rank the top 5 most relevant FAQs based on:
16 - Semantic similarity to the user’s intent
17 - Specificity to the question
18 - Likelihood of containing the information the user needs
194. Provide a confidence score (0-100) for each match
205. Explain your reasoning process
21For banking-related queries, consider:
22- Security concerns take priority
23- Account access questions require specific authentication-related FAQs
24- Transaction questions should match to relevant transaction FAQs
25- General inquiries should match to general information FAQs
26You must produce your analysis as a JSON object according to the following schema:
27{
28"user_utterance": "The original user utterance",
29"intent_analysis": "A thorough analysis of what the user is asking for, including likely underlying needs",
30"primary_banking_category": "The main banking category this query falls under (Account Management, Security, Transactions, etc.)",
31"relevant_faqs": [
32{
33"faq": "The title of the FAQ",
34"relevance_score": 95,
35"reasoning": "Detailed explanation of why this FAQ is relevant to the user’s query"
36},
37...
38],
39"confidence_in_mapping": "HIGH/MEDIUM/LOW",
40"explanation_of_confidence": "Brief explanation of why you’re confident or uncertain about these mappings",
41"recommended_clarification_question": "If confidence is MEDIUM or LOW, provide a question that would help clarify the user’s intent"
42}
43If there are fewer than 5 relevant FAQs, only include those that are actually relevant.
44Available FAQs:
45[List of FAQs...]
46User Utterance: "{user_utterance}"
47"""’
Listing 2: FAQ Mapper Without Embeddings Prompt
A.2.2 FAQ mapper with answers
Both the FAQ mapper with answers agents follow a similar structure but includes answer content:
1system_message = """
2You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
3IMPORTANT GUIDELINES:
41. Analyze the user’s intent thoroughly
52. Match the intent to the most relevant FAQs based on both the FAQ question and its answer content
63. Rank FAQs by relevance (0-100 scale)
74. Provide clear reasoning for each match, considering the full context of the FAQ including its answer
85. Return exactly 5 FAQs unless there are fewer relevant ones
96. Be precise - banking customers need accurate information
10"""
11user_message = """
12You will be given a user utterance and a list of available FAQs with their answers. Your task is to:
131. Analyze what the user is truly asking about (identify the core intent)
142. Search through the available FAQs for relevant matches
153. Rank the top 5 most relevant FAQs based on:
16 - Semantic similarity to the user’s intent
17 - Specificity to the question
18 - Whether the answer content directly addresses the user’s needs
19 - Likelihood of containing the information the user needs
204. Provide a confidence score (0-100) for each match
215. Explain your reasoning process
22For banking-related queries, consider:
23- Security concerns take priority
24- Account access questions require specific authentication-related FAQs
25- Transaction questions should match to relevant transaction FAQs
26- General inquiries should match to general information FAQs
27You must produce your analysis as a JSON object according to the following schema:
28{
29"user_utterance": "The original user utterance",
30"intent_analysis": "A thorough analysis of what the user is asking for, including likely underlying needs",
31"primary_banking_category": "The main banking category this query falls under (Account Management, Security, Transactions, etc.)",
32"relevant_faqs": [
33{
34"faq": "The title of the FAQ",
35"relevance_score": 95,
36"reasoning": "Detailed explanation of why this FAQ is relevant to the user’s query, including how the answer content addresses the query"
37},
38...
39],
40"confidence_in_mapping": "HIGH/MEDIUM/LOW",
41"explanation_of_confidence": "Brief explanation of why you’re confident or uncertain about these mappings",
42"recommended_clarification_question": "If confidence is MEDIUM or LOW, provide a question that would help clarify the user’s intent"
43}
44Available FAQs with their Answers:
45[List of FAQs with answers...]
46User Utterance: "{user_utterance}"
47"""’
Listing 3: FAQ Mapper With Answers Prompt
A.2.3 ARQ vs. Chain-of-Thought implementation
For comparison, our CoT implementation uses a similar prompt but without the structured JSON format:
1system_message = """
2You are an expert FAQ annotation system for our banking application. Your role is to accurately map user utterances to the most relevant FAQs from the bank’s knowledge base.
3"""
4user_message = """
5You will be given a user utterance and a list of available FAQs. Your task is to find the most relevant FAQs for the user’s query.
6First, think step-by-step about what the user is asking for. Analyze their intent carefully.
7Then, identify which FAQs in the provided list would best address their query.
8Rank the top 5 most relevant FAQs, providing your reasoning for each selection.
9Available FAQs:
10[List of FAQs...]
11User Utterance: "{user_utterance}"
12"""
Listing 4: Chain-of-Thought Prompt Example
The crucial difference between our ARQ and CoT implementations is that ARQ provides explicit queries in a structured format that guide the LLM through a systematic reasoning process, while CoT allows for more free-form reasoning without explicit structure. This structured approach helps maintain consistency and reliability in the agents’ reasoning, particularly for complex banking queries.
A.3 Error analysis
We analyzed cases where our framework failed to rank the correct FAQ first. The most common error types were:
  • Ambiguous queries: Queries that could be interpreted in multiple ways, e.g., "fash" (fashion vs. phishing).
  • Technical terminology: Queries using technical banking terms that were not well-represented in the training data.
These findings suggest areas for future improvement, such as better handling of ambiguous queries.
A.4 Monitoring and evaluation
Continuous monitoring and evaluation are essential for maintaining system performance. We recommend:
  • Implementing a feedback loop where user interactions are logged and analyzed
  • Regularly retraining the system with new examples
  • Conducting periodic human evaluations of system outputs
  • Setting up automated monitoring for anomalies or performance degradation
A.5 Public dataset details
A.5.1 LCQMC dataset
The Large-scale Chinese Question Matching Corpus (LCQMC) Liu et al. [2018] is designed for the task of question matching, which aims to determine whether two questions have the same intent. The dataset contains over 260,000 question pairs collected from Baidu Knows, a popular Chinese community question-answering service.
For our experiments, we adapted this dataset to our FAQ annotation task by:
  • Treating one question in each pair as a "user utterance"
  • Treating the other question as an "FAQ question"
  • Using the provided binary label (0 or 1) to determine relevance
  • Creating a pool of potential FAQ questions from the dataset
This adaptation allowed us to evaluate our framework on a large-scale question matching task in a different language, demonstrating its cross-lingual capabilities.
A.5.2 FiQA dataset
The FiQA dataset is a financial domain question answering dataset created for the Financial Question Answering Task of the WWW’18 conference. It contains user questions about financial topics along with expert-provided answers.
For our experiments, we adapted this dataset to our FAQ annotation task by:
  • Using the original questions as FAQ questions
  • Creating variations of these questions as user utterances
  • Ensuring that the variations maintain the same intent as the original questions
  • Using the associated answers as FAQ answers
This adaptation allowed us to evaluate our framework on a finance-specific dataset that is different from our internal banking dataset, providing insights into the framework’s domain generalization capabilities.
A.6 Judge agent analysis
We analyzed the judge agent’s reasoning to understand how it makes decisions. Table 8 shows examples of reasonings provided by the judge agent.
Table 8: Examples of judge agent reasoning
Utterance Top FAQ Reasoning for Top FAQ
Lost deb Lock and unlock your credit and debit cards This FAQ is highly relevant as it provides information on securing a lost debit card by locking it, which is a primary concern when a card is lost.
sba What about my business accounts? This FAQ directly addresses business accounts, which are highly relevant to SBA-related inquiries. It provides information on deposit insurance for business accounts, which is pertinent to small business owners.
How much money do i have What is account preview? This FAQ directly explains how users can view limited account information, such as balances, before signing into the app, which is exactly what the user is asking about.
The judge agent consistently provides detailed reasoning that considers both the semantic similarity and the practical relevance of each FAQ. This not only improves accuracy but also makes the system more interpretable and trustworthy.
A.7 Practical considerations
Deploying a multi-agent FAQ annotation system in a production banking environment requires careful consideration of several practical factors:
A.7.1 Latency and resource requirements
Our framework involves multiple LLM calls, which can introduce latency. Table 9 shows the average latency for different configurations.
Table 9: Latency Analysis
Configuration Avg. Latency (ms)
Single Agent (Direct) 450
Single Agent (w/ Embedding) 550
MAFA (Sequential) 2,800
MAFA (Parallel) 650
By parallelizing agent calls and implementing efficient caching strategies, we can reduce latency to acceptable levels for production use. Our parallel implementation reduces the overall latency from 2,800 ms to 650 ms, making it comparable to a single agent approach while providing significantly better accuracy.
A.7.2 Adapting to new domains
Our experiments on public benchmarks demonstrate that the MAFA framework can be adapted to new domains with relatively minimal effort. Key considerations for domain adaptation include:
  • Few-shot example selection: Selecting appropriate few-shot examples for the new domain is crucial for performance.
  • Domain-specific instructions: Updating the agent prompts to include domain-specific terminology and considerations.
  • Answer integration: For domains where context from answers is important, ensuring that this information is properly integrated into the system.
Our findings suggest that the core architecture of MAFA is domain-agnostic, with the specialized few-shot example strategy providing a mechanism for efficient adaptation to new domains.
A.8 Societal Impacts
The paper primarily focuses on technical aspects without adequately addressing the broader societal implications of this research. Future revisions should include a comprehensive discussion of societal impacts. On the positive side, automated FAQ annotation systems could significantly enhance financial inclusion by making banking information more accessible to traditionally underserved populations, reduce customer frustration through faster and more accurate responses to multi-lingual queries, and enable financial institutions to scale their support services more efficiently across diverse language communities. The framework could also reduce the cognitive load on customer service representatives by handling routine inquiries, allowing them to focus on more complex customer needs requiring human empathy and judgment.
However, these benefits must be weighed against potential concerns. The progressive automation of customer service and annotation functions could accelerate workforce displacement, particularly affecting entry-level positions that have traditionally served as career entry points. This may disproportionately impact certain demographic groups overrepresented in these roles. Additionally, automated systems might perpetuate or amplify existing biases in training data, potentially leading to inequality in service quality across different customer segments. There are also privacy considerations regarding how user query data is stored, processed, and potentially repurposed when developing and refining these systems.
Future work should explore responsible deployment strategies, including retraining programs for affected workers, ongoing bias monitoring frameworks, and transparent disclosure to customers about when they are interacting with automated systems versus human representatives. The research community should also consider developing metrics that evaluate not just technical performance but also fairness, accessibility, and overall societal benefit when designing and implementing such systems.